# A Class of Countably Covered Two-Dimensional Sofic Shifts

### v1

@@ 4 Construction toolbox
In this section we present several auxiliary constructions of countable SFTs that are used as part of the main construction. Many of them were used in some form in [10], but we restate them here both for completeness and for ease of reuse.
Our most fundamental building block is the grid shift, a standard example of a nontrivial two-dimensional SFT. In our constructions it acts as a way of transferring information from one part of a configuration to everywhere else, as a scaffold on which we can implement computation, and as a “clock” that periodically resets said computation.
@@
The following is essentially [10, Example 14]. It allows us to constrain the number of nonzero columns in a configuration of $(A^{\mathbb{Z}})^{\dagger}$ by the size of a grid on a separate layer.
@@
Fix a finite alphabet $A$ containing 00. There exists an SFT $X_{\mathrm{count}}\subset(A^{\mathbb{Z}})^{\dagger}\times X_{\mathrm{grid}}\times A_{\mathrm{count}}^{\mathbb{Z}^{2}}$ with the following properties.
@@
    For each $n\in\mathbb{N}$, the set $X_{\mathrm{count}}(n):=\left\{z\in X_{\mathrm{count}}\;\middle|\;\pi_{\mathrm{grid}}(z)=x_{\mathrm{grid}}(n)\right\}$ is countable, and the projection $\pi_{A}(X_{\mathrm{count}}(n))$ is exactly the set of configurations of $(A^{\mathbb{Z}})^{\dagger}$ with at most $n-2$ nonzero columns.
@@
Note that we could enforce a bound of $n$ nonzero columns in item 1 instead of $n-2$, but this would require a slightly more complex tile set.
@@
The alphabet $A_{\mathrm{count}}$ consists of two layers of Wang tiles, which are shown in Figure 2. The first layer is the counter layer, and the second one is the synchronization layer. The alphabet of $X_{\mathrm{count}}$ also contains an $A$-layer and a grid layer. Each layer must respect the local matching rules of its Wang tiles. In addition, the layers are constrained by the following rules, where $c\in A\times A_{\mathrm{grid}}\times A_{\mathrm{count}}$ is an arbitrary symbol:
@@
    The grid layer of $c$ contains a horizontal line if and only if its counting layer contains a solid horizontal line (a tile from group 3 in Figure 2). In this case, the counting layer contains a white dot if and only if the synchronization layer contains a dashed vertical line.
@@
    If the counting layer of $c$ contains a solid vertical line (a tile from group 2 in Figure 2), then the grid layer must contain a vertical line.
@@
    The grid layer of $c$ contains both a vertical and a horizontal line if and only if the counting layer is in both groups 2 and 3.
@@
    If the counting layer of $c$ has a dashed horizontal line on its west border (so the tile is in group 1 of Figure 2), then the line turns north if and only if the $A$-layer is nonzero.
We describe the structure of configurations $z\in X_{\mathrm{count}}$; see Figure 3 for reference. Suppose that the grid layer of $z$ contains a grid of finite side length $n$, that is, $z\in X_{\mathrm{count}}(n)$. By rules 1 and 2, this grid completely determines the positions of all solid horizontal lines of the counting layer and all solid vertical lines of the synchronization layer. By rule 4 it also determines the positions of solid vertical lines of the counting layer, which must coincide with those of the other layers. The bi-infinite stripe between each pair of solid horizontal lines of the counting layer contains a single bi-infinite dashed path. By rule 5 the path goes up by one tile whenever it meets a nonzero column on the $A$-layer, and continues horizontally otherwise. Since there are $n-1$ rows between two adjacent grid lines, the $A$-layer of $z$ can contain at most $n-2$ nonzero columns.
The positions of the dashed paths of different stripes in $z$ are synchronized by the diagonal lines of the counting layer: when the path crosses a solid vertical line, a diagonal line is sent to the southeast, which eventually meets a solid horizontal line and produces a white dot. By rule 1 the positions of the dots must match the intersections of the dashed lines of the synchronization layer and the horizontal lines of the grid layer. Thus there are a finite number of options for the positions of the dashed paths, each of which determines the synchronization layer. Since there are also finitely many choices for the $A$-layer, the set $X_{\mathrm{count}}(n)$ is countable. We also see that the $A$-layer can be chosen freely, as long as it has at most $n-2$ nonzero columns. This proves item 1.
For item 2, fix a configuration $x\in(A^{\mathbb{Z}})^{\dagger}$, and let $z\in X_{\mathrm{count}}$ be such that $\pi_{A}(z)=x$. We show that there are a countable number of choices for $z$. If the grid layer of $z$ contains a grid of finite side length, the claim follows from item 1 Suppose it does not contain such a grid. Then the counting layer contains at most one solid horizontal line and at most one solid vertical line, and the synchronization layer contains at most one solid vertical line. In particular, there are countably many choices for the contents of the synchronization layer. Since there must be a solid horizontal line between each pair of dashed paths on the counting layer, $z$ contains at most two such paths. There are a countable number of choices for the number of these paths and the positions at which they cross the y-axis. Since we fixed the $A$-layer $\pi_{A}(z)=x$, after these choices the dashed paths are completely determined. Finally, between two diagonal lines of the counting layer there must be a thick horizontal line, so the number of choices for them is likewise countable. ∎
(a)123(b)
[…]
@@
Figure 3: A sample configuration of $X_{\mathrm{count}}$. The counting layer is shown, as are the dashed lines of the synchronization layer. The two arrows mark nonzero columns of the $A$-layer.
@@
We describe a way of simulating computation by nondeterministic oracle counter machines in SFTs. We do not explicitly define a $k$-counter machine, but it consists of a finite oracle alphabet $A$, a finite state set $Q$, and initial state $q_{0}\in Q$, and a transition relation $\delta$. The machine has $k$ counters that store one natural number each, which are initialized to 0, and possibly an infinite oracle tape $x\in A^{\mathbb{N}}$. If the oracle tape exists, then one of the counters is designated as the oracle counter. Alternatively, the oracle tape may be a bi-directional configuration $x\in A^{\mathbb{Z}}$, in which case each counter will store an integer instead of a natural number. In this latter case the counter machine is called bidirectional.
In one computation step the machine will:
  • check the relative order of the counters, and which of them have value 0,
  • check the symbol $x_{n}$, where $n$ is the current value of the oracle counter,
  • based on the above data, the current state, and possibly a nondeterministic choice, change the value of any counters by at most 1 and assume a new state.
@@
Let $M$ be a one-directional $k$-counter machine as above. Define $H_{M}=H^{\prime}_{M}\times\{{\leftarrow},{\rightarrow}\}$ and $A_{M}=\{\#\}\cup((H_{Q}\cup\{L,R\})\times\{Z,P\}^{k}\times A)$, where $H^{\prime}_{M}$ is an auxiliary alphabet. Define an SFT $X_{M}\subset A_{M}^{\mathbb{Z}^{2}}$ as follows. If a horizontal row of $x\in X$ contains a $-\ell$ segment of non-##\## symbols starting at $\vec{v}$, then a length-$(\ell-1)$ segment must start below it at $\vec{v}-(0,1)$ and a length-$(\ell+1)$ segment above it at $\vec{v}+(0,1)$. We interpret time as increasing upward and introduce (nondeterministic) rules that transform a segment into the one above it.
@@
Each segment consists of $k+1$ layers. Its first layer must be in the language $L^{*}H_{M}R^{*}$ (storing the position, direction and current transition in a “zig-zag head”) and the next $k$ layers must be in $P^{*}Z^{*}$ (storing the value of one counter). The final layer is in $A^{*}$, and each segment is a prefix of the one above; their limit corresponds to the oracle tape. The head travels along the segment in the direction of its arrow component, with speed 2 to the east and speed 1 to the west. When it steps east over a counter, it updates the counter’s value based on its $H^{\prime}_{M}$-component; when it steps west over one, it records its place in the ordering of the counters, as well as the symbol on the $A$-layer in case of the oracle counter. When it reaches the east border of the segment, it turns west, and upon reaching the west border, it updates its $H^{\prime}_{M}$-component based on the information it stores. In this way, one west-east sweep corresponds to one computation step of $M$.
@@
If the machine is bidirectional, we modify the computation cone to grow both to the left and to the right. The left and right halves of the cone have different background colors, and the central column of the cone corresponds to counter value 00.
@@
The most important property of $X_{M}$ is that if we ignore the $A$-component, the set of “degenerate” configurations – those that do not contain a simulated computation of $M$ – is countable. Indeed, such a configuration has at most one connected region of non-##\## cells, which is shaped like a half-plane and contains at most one infinite sweep of the zig-zag head. We will modify this basic construction as needed, in particular by allowing the computation cones to be cut short when $M$ halts, and attaching them to grid cells.
@@
Recall from [6, Section 11.2] that any Turing machine computation can be simulated by a counter machine computation in a way that incurs an exponential blowup in the number of computation steps. In our construction, the first $n$ simulated steps of a counter machine computation fit in an exponential-size square pattern. Hence, the blowup from a Turing machine computation to a pattern is doubly exponential.
@@
Take the grid shift $X_{\mathrm{grid}}$ and superimpose on each gray region a computation of a counter machine using a variation of the SFT $X_{M}$ in which the computation cones are allowed to be cut off at any point after the simulated machine has halted, and the oracle tape is blank. We give $M$ one counter that corresponds to the value of $n$, and four counters that correspond to the four cardinal directions. The machine $M$ first nondeterministically sets each of these counters to some value (by incrementing them some number of times) and never modifies them again. Then it simply checks whether the 5-tuple $(n,t_{e},t_{n},t_{w},t_{s})$ they form satisfies $(t_{e},t_{n},t_{w},t_{s})\in T_{n}$, halting if it does.
We also use signals to transfer information between the machines in adjacent grid cells: we enforce each simulated machine in a single configuration to choose the same value of $n$, and for the direction counters to be related as in a Wang tiling (so the north counter of a grid cell must equal the south counter of its north neighbor, and analogously for west and east). See [9, Theorem 6.2] for a more detailed construction and analysis.
@@
We describe a modification to the SFT $X$ of Example 14 that adds a vertically constant layer over a separate alphabet $A$ and gives the simulated tile sets access to it. For this, we assume that each tile set $T_{n}$ has a separate layer over the auxiliary alphabet $B=\{0,1,2\}\cup(A\times\{0,1\})$, and all $2\times 1$ patterns except $00$, $0a$, $a1$, $a2$, $11$, $12$ and $22$ for $a\in A\times\{0,1\}$ are forbidden on this layer. This layer is used to “probe” the contents of the $A$-layer.
We construct the counter machine $M$ so that it contains a special probe state $q_{b}$ for each $b\in B$. It will enter a probe state $q_{b}$ at some point during its computation if and only if the $B$-layer of the tile it simulates equals $b$.
We superimpose on $X\times(A^{\mathbb{Z}})^{\dagger}$ a configuration over the tile set $A_{\mathrm{probe}}$ of Figure 4, called the probe layer, and further constrain it using the following rules:
@@
    The thick horizontal and vertical lines of the probe layer must coincide with those of the grid layer.
@@
    The highlighted tiles in Figure 4, which include only those with $b^{\prime}\in A\times\{0\}$, are called the probe tiles. The $A$-part of the symbol $b^{\prime}$, $a_{0}$ or $a_{1}$ of a probe tile must be equal to the symbol on the $A$-layer.
@@
Denote by $H\subseteq X\times(A^{\mathbb{Z}})^{\dagger}\times A_{\mathrm{probe}}^{\mathbb{Z}^{2}}$ the SFT thus obtained. The probe layer of a sample configuration is depicted in Figure 5.
Consider a configuration $(x,y,z)\in H$, where $x\in X$ contains a grid of width $m\geq 0$ that simulates a configuration of tile set $T_{n}$ for some $n\in\mathbb{N}$. By item 1, the horizontal and vertical lines of the probe layer $z$ are aligned with the grid, and by item 3 and our assumption on $M$, each vertical line segment in $z$ holds the value of the $B$-layer of the tile of $T_{n}$ simulated on its right. By our assumption on $T_{n}$, each row of the simulated tiling contains at most one symbol from $A\times\{0,1\}$, and on its right some number of $1$-symbols, all other symbols coming from $\{0,2\}$. Thus, each row of the grid contains at most one vertical segment $L$ holding a value $(a,c)\in A\times\{0,1\}$ on the probe layer.
If the next $k>0$ segments to the right hold $1$-symbols, the rightmost one of these emits a dashed line to the left on the southmost row of the grid cell. This signal is raised to the north by one step each time it crosses a segment holding a $1$-symbol, and upon collision with $L$, sends a diagonal signal to the southeast. This signal carries the same symbol $(a,c)$, and as it hits the south border of the grid cell at distance $k$ from segment $L$, rule 2 guarantees that the value of the $A$-layer at that position is $a$. We also see that $k<m$. Because of the black disks at the borders of certain tiles, $c=1$ if and only if the diagonal signal hits the tile neighboring the next segment to the right from $L$, if and only if $k=m-1$. In the case $k=0$, by rule 2 the segment $L$ holds the symbol $(a,0)$, where $a\in A$ is the symbol of the $A$-layer on the same column. In this way, the $B$-layer of the simulated tilings can be used to access the width of the grid layer and the full contents of the $A$-layer.
As for the cardinality of the new SFT, note first that if the grid layer contains a finite-width grid, then the probe layer is completely determined by the other layers. Conversely, if there is no finite-width grid, then the probe layer too contains no parallel thick lines, and a simple case analysis shows that the number of valid contents for the probe layer is countable when the other layers are fixed.
@@
Figure 4: The tile set $A_{\mathrm{probe}}$. Here $b$ ranges over $B$, $b^{\prime}$ ranges over $\{0,1,2\}\cup(A\times\{0\})$, $a$ ranges over $A$, and $a_{i}$ for $i\in\{0,1\}$ ranges over $A\times\{i\}$.

### v2

@@ 4 Construction toolbox
In this section we present several auxiliary constructions of countable SFTs that are used to prove Theorems 6 and 8. Many of them were first used in some form in [12], but we restate them here both for completeness and for ease of reuse.
Our most fundamental building block is the grid shift, a standard example of a nontrivial two-dimensional SFT. In our constructions it acts as a way of transferring information from one part of a configuration to everywhere else, as a scaffold on which we can implement computation and various geometric gadgets, and as a “clock” that periodically resets said computation.
@@
The following is essentially Example 14 of [12]. It allows us to constrain the number of nonzero columns in a configuration of $(A^{\mathbb{Z}})^{\dagger}$ by the size of a grid on a separate layer.
@@
Fix a finite alphabet $A$ containing 0. There exists an SFT $X_{\mathrm{count}}\subset(A^{\mathbb{Z}})^{\dagger}\times X_{\mathrm{grid}}\times A_{\mathrm{count}}^{\mathbb{Z}^{2}}$ with the following properties.
@@
    For each $n\in\mathbb{N}$, the set $X_{\mathrm{count}}(n):=\left\{z\in X_{\mathrm{count}}\;\middle|\;\pi_{\mathrm{grid}}(z)=x_{\mathrm{grid}}(n)\right\}$ is countable, and the projection $\pi_{A}(X_{\mathrm{count}}(n))$ is exactly the set of configurations of $(A^{\mathbb{Z}})^{\dagger}$ with at most n2n-2 nonzero columns.
@@
Note that we could enforce a bound of $n$ nonzero columns in item 1 instead of n2n-2, but this would require a slightly more complex tile set.
@@
The alphabet $A_{\mathrm{count}}$ consists of two layers of Wang tiles, which are shown in Figure 2. The first layer is the counting layer, and the second one is the synchronization layer. The alphabet of $X_{\mathrm{count}}$ also contains an $A$-layer and a grid layer.
(a)(b)
The counting layer forms a unary counter in each vertical segment between two horizontal grid lines. This counter is incremented whenever it crosses a nonzero column on the $A$-layer. It cannot increase beyond the grid lines, which restricts the number of nonzero columns exactly as desired. The counter may have a nonzero initial value, so if we could choose it independently between any two horizontal grid lines, our SFT would be uncountable. Thus, the synchronization layer is used to force each counter to have the same initial value.
Each layer must respect the local matching rules of its Wang tiles. In addition, the layers are constrained by the following rules, where $c\in A\times A_{\mathrm{grid}}\times A_{\mathrm{count}}$ is an arbitrary symbol:
@@
    The grid layer of $c$ contains a horizontal line if and only if its counting layer contains a thick horizontal line (with a shaded region above it).
@@
    If the synchronization layer of $c$ contains a horizontal line, then the grid layer contains a horizontal line.
@@
    If the synchronization layer of $c$ contains a white dot, then the counting layer of $c$ has a thin horizontal line on its west border (with a shaded region below it).
@@
    If the counting layer of $c$ has a thin horizontal line on its west border, then the line turns north if and only if the $A$-layer is nonzero.
$\downarrow\downarrow$
@@
Figure 3: A sample configuration of $X_{\mathrm{count}}$. Depicted are the grid lines of the grid layer, the counting layer, and the synchronization layer. The two arrows mark nonzero columns of the $A$-layer.
We describe the structure of configurations $z\in X_{\mathrm{count}}$; see Figure 3 for reference. Suppose that the grid layer of $z$ contains a grid of finite side length $n$, that is, some translate of $z$ is in $X_{\mathrm{count}}(n)$. By rules 1 and 2, this grid completely determines the positions of all thick horizontal lines of the counting layer and all solid vertical lines of the synchronization layer. By rule 3 it also determines the positions of horizontal lines of the synchronization layer, which must coincide with those of the other layers. The bi-infinite stripe between each pair of thick horizontal lines of the counting layer contains a single bi-infinite thin path. By rule 5 the path goes up by one tile whenever it meets a nonzero column on the $A$-layer, and continues horizontally otherwise. Since there are n1n-1 rows between two adjacent grid lines, the $A$-layer of $z$ can contain at most n2n-2 nonzero columns.
The positions of the thin paths of different stripes in $z$ are synchronized by the synchronization layer. Between each pair of vertical grid lines lies a single dashed vertical line of the synchronization layer. Wherever it crosses a horizontal grid line, a diagonal line is sent to the northwest. By rule 4, it must meet the path of the counting layer and a vertical grid line at the same position (at a white dot). Thus, the heights of the counters are synchronized at each vertical grid line. Since there are also finitely many choices for the $A$-layer, the set $X_{\mathrm{count}}(n)$ is countable. We also see that the $A$-layer can be chosen freely, as long as it has at most n2n-2 nonzero columns. This proves item 1.
For item 2, fix a configuration $x\in(A^{\mathbb{Z}})^{\dagger}$, and let $z\in X_{\mathrm{count}}$ be such that $\pi_{A}(z)=x$. We show that there are a countable number of choices for $z$. If the grid layer of $z$ contains a grid of finite side length, the claim follows from item 1.
Suppose it does not contain such a grid. Then the counting layer contains at most one thick horizontal line, and hence at most two thin paths. Since the $A$-layer is fixed, the position at which a thin path crosses the y-axis determines the entire path, so the number of choices for the counting layer is countable. The synchronization layer contains at most one solid vertical line (and hence at most two dashed vertical lines), and its horizontal line segments (the south sides of the dark gray triangles) are constrained to a single line, which is the sole horizontal line of the grid layer if one exists. There cannot be more than two infinite diagonal lines, and the positions of finite diagonal segments are determined by the other lines. Thus, we have countably many choices for the synchronization layer as well. ∎
We now introduce a method for simulating computation in countable SFTs. To our best knowledge, its first use is in [11].
@@
We describe a way of simulating computation by nondeterministic oracle counter machines in SFTs. We do not explicitly define a $k$-counter machine, but it consists of a finite oracle alphabet $A$, a finite state set $Q$, an initial state $q_{0}\in Q$, and a transition relation $\delta$. The machine has $k$ counters that store one natural number each, which are initialized to 0, and possibly an infinite oracle tape $x\in A^{\mathbb{N}}$. If the oracle tape exists, then one of the counters is designated as the oracle counter. Alternatively, the oracle tape may be a bi-directional configuration $x\in A^{\mathbb{Z}}$, in which case each counter will store an integer instead of a natural number. In this latter case the counter machine is called bidirectional.
In one computation step the machine will check the relative order of its counters, which of them have value 0, as well as the symbol $x_{n}\in A$, where $n$ is the current value of the oracle counter. Based on the above data and the current state, it will change the value of any of the counters by at most 1 and assume a new state. The validity of a given transition is determined by the relation $\delta$. If there are two or more valid actions, one is chosen nondeterministically.
@@
Let $M$ be a one-directional $k$-counter machine as above. Define $H_{M}=H^{\prime}_{M}\times\{{\leftarrow},{\rightarrow}\}$ and $A_{M}=\{\#\}\cup((H_{M}\cup\{L,R\})\times\{Z,P\}^{k}\times A)$, where $H^{\prime}_{M}$ is an auxiliary alphabet that contains a transition from $\delta$, a partial ordering of the $k$ counters, and a symbol of $A$. Define an SFT $X_{M}\subset A_{M}^{\mathbb{Z}^{2}}$ as follows. First, for each position $\vec{v}\in\mathbb{Z}^{2}$, there is a non-#\# symbol at $\vec{v}$ if and only if $\vec{v}+(0,1)$ and $\vec{v}+(1,1)$ both contain non-#\# symbols. Then each connected component of non-#\# cells is shaped like a translate of $\{(i,j)\mid j\geq 0,0\leq i\leq j\}$. Such a component is called a computation cone, and its southmost cell is called its vertex.
We interpret time as increasing upward and introduce (nondeterministic) local rules that transform a horizontal row of a computation cone into the one above it. Each row will contain exactly one symbol $(q,d)\in H_{M}$ on the first layer, interpreted as a “zig-zag head” traveling in the direction $d\in\{{\leftarrow},{\rightarrow}\}$. The symbols $L$ and $R$ fill the rest of the row on this layer to the left and right of the head, respectively. The zig-zag head bounces between the east and west borders of the computation cone, and one such back-and-forth sweep corresponds to a single computation step of $M$. Each counter is represented by the number of $P$-symbols on its own layer, which has the form $P^{m}Z^{n}$ for some $m,n\geq 0$. As the zig-zag head travels the length of the computation cone, it records the relative order of the counter values in its internal state, and updates counter values based on the transition it contains.
Let us go into a little more detail. Each row of a computation cone consists of $k+2$ layers. The first layer must be in the language $L^{*}H_{M}R^{*}$ (storing the position and internal state of the zig-zag head) and the next $k$ layers must be in $P^{*}Z^{*}$ (storing the value $m\geq 0$ of one counter as $P^{m}Z^{n}$). The final layer is in $A^{*}$, and this layer of each row of the cone is a prefix of the one above; their limit corresponds to the oracle tape. At the vertex of the cone, the head contains an arbitrary transition from the initial state $q_{0}$ to some other state.
@@
The head travels in the direction of its arrow component, with speed 2 to the east and speed 1 to the west. The simulation of one computation step of $M$ begins at the west border of the cone, where the zig-zag head picks a new transition $t\in\delta$ into its $H^{\prime}_{M}$-component and walks east. When it steps east over a $PZ$-pattern on one of the counter layers, it updates that counter’s value if specified to do so by the transition $t$. Upon reaching the east border of the cone, it turns west and resets the partial order in its state into an emtpy one. When the head steps west over a $PZ$-pattern, it records into its internal state the place of the counter in the ordering of all counters, as well as the symbol on the $A$-layer in case of the oracle counter. Upon returning to the west border, the head picks a new valid transition from $\delta$ (based on the $A$-symbol and the now total order it stores) and begins the next computation step.
@@
If the machine is bidirectional, we modify the computation cone to grow both to the left and to the right. The left and right halves of the cone have different background colors, and the central column of the cone corresponds to counter value 0.
@@
The most important property of $X_{M}$ is that if we ignore the $A$-component, the set of “degenerate” configurations – those that do not contain a simulated computation of $M$ – is countable. Indeed, such a configuration has at most one connected region of non-#\# cells, which is shaped like a half-plane and contains at most one infinite sweep of the zig-zag head. We will modify this basic construction as needed, in particular by allowing the computation cones to be cut short when $M$ halts, and attaching them to grid cells.
@@
Recall from [8, Section 11.2] that any Turing machine computation can be simulated by a counter machine computation in a way that incurs an exponential blowup in the number of computation steps. In our construction, the first $n$ simulated steps of a counter machine computation fit in an exponential-size square pattern. Hence, the blowup from a Turing machine computation to a pattern is doubly exponential.
@@
Take the grid shift $X_{\mathrm{grid}}$ and superimpose on each gray region a computation of a counter machine using a variation of the SFT $X_{M}$ in which the computation cones are allowed to be cut off at a horizontal border of the grid provided that the simulated machine has halted, and the oracle tape is blank. We give $M$ one counter that corresponds to a choice of which tile set $T_{n}$ to use, and four counters that correspond to the four sides of a tile. The machine $M$ first nondeterministically sets each of these counters to some value (by incrementing them some number of times) and never modifies them again. Then it simply checks whether the 5-tuple $(n,t_{e},t_{n},t_{w},t_{s})$ they form satisfies $(t_{e},t_{n},t_{w},t_{s})\in T_{n}$, halting if it does. In this case we say that the machine, or the grid cell containing it, simulates the tile $(t_{e},t_{n},t_{w},t_{s})$ of $T_{n}$.
We also use signals to transfer information between the machines in adjacent grid cells. We enforce each simulated machine in a single configuration to choose the same value of $n$, and for the direction counters to be related as in a Wang tiling (so the north counter of a grid cell must equal the south counter of its north neighbor, and analogously for west and east). In this way, the grid cells of a valid configuration of the SFT together simulate a configuration of $T_{n}$.
The width of the grid is not constrained, except that if a simulated machine does not have enough space to finish its computation, a tiling error results. However, for each $n$ there exists $k\geq 1$ such that any grid of width at least $k$ can simulate an arbitrary configuration of $T_{n}$.
See [11, Theorem 6.2] for a more detailed construction and analysis.
Consider the following scenario. We have a vertically constant layer over $A$ that we wish to constrain. The constraints we wish to apply depend on some integer value $n\geq 1$, which in our case is the minimum distance between two nonzero symbols. We can add a grid layer, somehow extract $n$ from the $A$-layer of a nondegenerate configuration and constrain the grid cells to have width $n$. Also, from $n$ we can compute a Wang tile set $T_{n}$ that, when coupled with the $A$-layer in a suitable way, implements the constraints we want. We would like to add a layer given by Example 17 that simulates all the tile sets $T_{n}$, and couple it with both the grid layer (where it gets the value $n$) and the $A$-layer (which it constrains). The issue is that the simulated configurations of $T_{n}$ are not on the same “scale” as the $A$-layer: they are stretched over grids whose width we cannot control, so that each simulated tile of $T_{n}$ contains an $m\times m$ pattern of $A$-symbols, for some unknown and typically very large $m\geq 1$.
In the next example, which a new contribution, we provide an interface through which the simulated tilings of $T_{n}$ can still access and constrain arbitrary columns of the $A$-layer. The tile sets $T_{n}$ have to be specifically designed to make use of this interface; in the above scenario, we would probably have to redesign them completely.
@@
We describe a modification to the SFT $X$ of Example 17 that adds a vertically constant layer over a separate alphabet $A$, called the base alphabet layer, and gives the simulated tile sets access to it. For this, we assume that each tile set $T_{n}$ has a separate layer, called the probe interface layer, over the auxiliary alphabet $B=\{0,1,2\}\cup(A\times\{0,1\})$. All $2\times 1$ patterns except $00$, $0a$, $a1$, $a2$, $11$, $12$ and $22$ for $a\in A\times\{0,1\}$ are forbidden on the probe interface layer. Hence, each row of this layer has the form $^{\infty}0a1^{k}2^{\infty}$ for some $a\in A\times\{0,1\}$ and $k\geq 0$ (or a degenerate version). As the name suggests, the probe interface layer is used to “probe” the contents of the base alphabet layer.
The interface between the base alphabet layer and the simulated tiling of $T_{n}$ is the following. As in Example 17, we have a grid layer, and each grid cell simulates a tile of $T_{n}$ for a common $n$. Suppose that on some row of the simulated tiling, simulated by $k+2$ consecutive grid cells, we have a word of the form $w=(a,c)1^{k}2$ on the probe interface layer, with $k\geq 0$ and $(a,c)\in A\times\{0,1\}$. Then $a$ must be equal to the symbol of the base alphabet layer on the $k$th column of the grid cell that simulates $(a,c)$. We think of this as the simulated tile set $T_{n}$ placing the word $w$ at a certain position of its probe interface layer in order to access a certain column of the non-simulated base alphabet layer. Also, $c=1$ if and only if the probed column is the rightmost one of the grid cell; this lets the tile set $T_{n}$ avoid attempting to probe beyond the grid cell that simulates $(a,c)$, which our construction does not allow. Note that this is not an essential restriction, since different rows of $T_{n}$ may probe columns of different grid cells, so that a single simulated configuration of $T_{n}$ may access every column of the base alphabet layer.
We construct the counter machine $M$ of Example 17 so that it contains a special probe state $q_{b}$ for each $b\in B$. It will enter a probe state $q_{b}$ at some point during its computation if and only if the probe layer of the tile it simulates equals $b$. Hence, during a finite computation $M$ will enter exactly one of the probe states before halting.
We superimpose on $X\times(A^{\mathbb{Z}})^{\dagger}$ a configuration over the tile set $A_{\mathrm{probe}}$ of Figure 4, called the probe implementation layer, and further constrain it using the following rules:
@@
    The thick horizontal and vertical lines of the probe implementation layer must coincide with those of the grid layer of $X$.
@@
    The highlighted tiles in Figure 4, which include only those with $b_{0}\in A\times\{0\}$, are called the probe tiles. The $A$-part of the symbol $b_{0}$, $a_{0}$ or $a_{1}$ of a probe tile must be equal to the symbol on the $A$-layer.
@@
Denote by $H\subseteq X\times(A^{\mathbb{Z}})^{\dagger}\times A_{\mathrm{probe}}^{\mathbb{Z}^{2}}$ the SFT thus obtained. By rule 1 we can think of the probe implementation layer as decorations over the grid that simulates copies of $M$. The probe implementation layer of a sample configuration is depicted in Figure 5, which the reader should consult to better understand the following explanation.
@@
Figure 4: The tile set $A_{\mathrm{probe}}$ of the probe implementation layer. Here $b$ ranges over $B$, $b_{0}$ ranges over $\{0,1,2\}\cup(A\times\{0\})\subset B$, $a_{*}$ ranges over $A\times\{0,1\}$, and $a_{i}$ for $i\in\{0,1\}$ ranges over $A\times\{i\}$.